# references.bib

### v2

@@ Abstract
There has been a recent surge in transformer-based architectures for learning on graphs, mainly motivated by attention as an effective learning mechanism and the desire to supersede handcrafted operators characteristic of message passing schemes. However, concerns over their empirical effectiveness, scalability, and complexity of the pre-processing steps have been raised, especially in relation to much simpler graph neural networks that typically perform on par with them across a wide range of benchmarks. To tackle these shortcomings, we consider graphs as sets of edges and propose a purely attention-based approach consisting of an encoder and an attention pooling mechanism. The encoder vertically interleaves masked and vanilla self-attention modules to learn an effective representations of edges, while allowing for tackling possible misspecifications in input graphs. Despite its simplicity, the approach outperforms fine-tuned message passing baselines and recently proposed transformer-based methods on more than 70 node and graph-level tasks, including challenging long-range benchmarks. Moreover, we demonstrate state-of-the-art performance across different tasks, ranging from molecular to vision graphs, and heterophilous node classification. The approach also outperforms graph neural networks and transformers in transfer learning settings, and scales much better than alternatives with a similar performance level or expressive power.
1 Introduction
@@
We investigate empirically the potential of a purely attention-based approach for learning effective representations of graph structured data. Typically, learning on graphs is modelled as message passing – an iterative process that relies on a message function to aggregate information from a given node’s neighbourhood and an update function to incorporate the encoded message into the output representation of the node. The resulting graph neural networks (GNNs) typically stack multiple such layers to learn node representations based on vertex rooted sub-trees, essentially mimicking the one-dimensional Weisfeiler–Lehman ($1$-WL) graph isomorphism test [wl79, wlneural24]. Variations of message passing have been applied effectively in different fields such as life sciences [STOKES2020688, Wong2023, gasteiger_dimenet_2020, Merchant2023, doi:10.1126/science.abj8754, doi:10.1126/science.add2187], electrical engineering [Chien2024], and weather prediction [doi:10.1126/science.adi2336].
@@
Despite the overall success and wide adoption of graph neural networks, several practical challenges have been identified over time. While the message passing framework is highly flexible, the design of new layers is a challenging research problem where improvements take years to achieve and often rely on hand-crafted operators. This is particularly the case for general purpose graph neural networks that do not exploit additional input modalities such as atomic coordinates. For instance, principal neighbourhood aggregation (PNA) is regarded as one of the most powerful message passing layers [10.5555/3495724.3496836], but it is built using a collection of manually selected neighbourhood aggregation functions, requires a dataset degree histogram which must be pre-computed prior to learning, and further uses manually selected degree scaling. The nature of message passing also imposes certain limitations which have shaped the majority of the literature. One of the most prominent examples is the readout function used to combine node-level features into a single graph-level representation, and which is required to be permutation invariant with respect to the node order. Thus, the default choice for graph neural networks and even graph transformers remains a simple, non-learnable function such as sum, mean, or max [buterez2022graph, Buterez2023, D3SC00841J]. Limitations of this approach have been identified by Wagstaff et al. [pmlr-v97-wagstaff19a], who have shown that simple readout functions might require complex item embedding functions that are difficult to learn using standard neural networks. Additionally, graph neural networks have shown limitations in terms of over-smoothing [gcnsmooth], linked to node representations becoming similar with increased depth, and over-squashing [alon2021on] due to information compression through bottleneck edges. The proposed solutions typically take the form of message regularisation schemes [godwin2022simple, Zhao2020PairNorm, cai2021graphnorm]. However, there is generally no consensus on the right architectural choices for building effective deep message passing neural networks. Transfer learning and strategies like pre-training and fine-tuning are also less ubiquitous in graph neural networks because of modest or ambiguous benefits, as opposed to large language models [Hu2020Strategies].
The attention mechanism [NIPS2017_3f5ee243] is one of the main sources of innovation within graph learning, either by directly incorporating attention within message passing [velickovic2018graph, brody2022how], by formulating graph learning as a language processing task [ying2021do, kim2022pure], or by combining vanilla GNN layers with attention layers [rampasek2022recipe, shirzad2023exphormer, buterez2022graph, Buterez2023]. However, several concerns have been raised regarding the performance, scalability, and complexity of such methods. Performance-wise, recent reports indicate that sophisticated graph transformers underperform compared to simple but tuned GNNs [tonshoff2023where, luo2024classicgnnsstrongbaselines]. This line of work highlights the importance of empirically evaluating new methods relative to strong baselines. Separately, recent graph transformers have focused on increasingly more complex helper mechanisms, such as computationally expensive pre-processing and learning steps [ying2021do], various different encodings (e.g., positional, structural, and relational) [ying2021do, kim2022pure, rampasek2022recipe, 10.5555/3618408.3619379], inclusion of virtual nodes and edges [ying2021do, shirzad2023exphormer], conversion of the problem to natural language processing [ying2021do, kim2022pure], and other non-trivial graph transformations [shirzad2023exphormer, 10.5555/3618408.3619379]. These complications can significantly increase the computational requirements, reducing the chance of being widely adopted and replacing GNNs.
Motivated by the effectiveness of attention as a learning mechanism and recent advances in efficient and exact attention, we introduce an end-to-end attention-based architecture for learning on graphs that is simple to implement, scalable, and achieves state-of-the-art results. The proposed architecture considers graphs as sets of edges, leveraging an encoder that interleaves masked and self-attention mechanisms to learn effective representations. The attention-based pooling component mimics the functionality of a readout function and it is responsible for aggregating the edge-level features into a permutation invariant graph-level representation. The masked attention mechanism allows for learning effective edge representations originating from the graph connectivity and the combination with self-attention layers vertically allows for expanding on this information while having a strong prior. Masking can, thus, be seen as leveraging specified relational information and its vertical combination with self-attention as a means to overcome possible misspecification of the input graph. The masking operator is injected into the pairwise attention weight matrix and allows only for attention between linked primitives. For a pair of edges, the connectivity translates to having a shared node between them. We focus primarily on learning through edge sets due to empirically high performance, and refer to our architecture as edge-set attention (ESA). We also demonstrate that the overall architecture is effective for propagating information across nodes through dedicated node-level benchmarks. The ESA architecture is general purpose, in the sense that it only relies on the graph structure and possibly node and edge features, and it is not restricted to any particular domain. Furthermore, ESA does not use positional, structural, relative, or similar encodings, it does not encode graph structures as tokens or other language (sequence) specific concepts, and it does not require any pre-computations.
Despite its apparent simplicity, ESA-based learning overwhelmingly outperforms strong and tuned GNN baselines and much more involved transformer-based models. Our evaluation is extensive, totalling $70$ datasets and benchmarks from different domains such as quantum mechanics, molecular docking, physical chemistry, biophysics, bioinformatics, computer vision, social networks, functional call graphs, and synthetic graphs. At the node level, we include both homophilous and heterophilous graph tasks, as well as shortest path problems, as these require modelling long-range interactions. Beyond supervised learning tasks, we explore the potential for transfer learning in the context of drug discovery and quantum mechanics [Buterez_Janet_Kiddle_Oglic_Liò_2024] and show that ESA is a viable transfer learning strategy compared to vanilla GNNs and graph transformers.
2 Related Work
An attention mechanism that mimics message passing and limits the attention computations to neighbouring nodes has been first proposed in GAT [velickovic2018graph]. We consider masking as an abstraction of the GAT attention operator that allows for building general purpose relational structures between items in a set (e.g., $k$-hop neighbourhoods or conformational masks extracted from $3$D molecular structures). The attention mechanism in GAT is implemented as a single linear projection matrix that does not explicitly distinguish between keys, queries, and values as in standard dot product attention and an additional linear layer after concatenating the representations of connected nodes, along with a non-linearity. This type of simplified attention has been labelled by subsequent work as static and was shown to have limited expressive power [brody2022how]. Brody et al. instead proposed dynamic attention, a simple reordering of operations in GAT, resulting in a more expressive GATv2 model [brody2022how] – however, at the price of doubling the parameter count and the corresponding memory consumption. A high-level overview of GAT in the context of masked attention is provided in Figure 1, along with the main differences to the proposed architecture. Similar to GAT, several adaptations of the original scaled dot product attention have been proposed for graphs [ijcai2021p214, dwivedi2021generalization], where the focus was on defining an attention mechanism constrained by the node connectivity and replacing the positional encodings of the original transformer model with more appropriate graph alternatives, such as Laplacian eigenvectors. These approaches, while interesting and forward-looking, did not convincingly outperform simple GNNs. Building up on this line of work, an architecture that can be seen as an instance of masked transformers has been proposed in SAN [kreuzer2021rethinking], illustrated in Figure 2. Attention coefficients there are defined as a convex combination of the scores (controlled by hyper-parameter $\gamma$) associated with the original graph and its complement. Min et al. [10.1145/3477495.3532031] have also considered a masking mechanism for standard transformers. However, the graph structure itself is not used directly in the masking process as the devised masks correspond to four types of prior interaction graphs (induced, similarity, cross neighbourhood, and complete sub-graphs), acting as an inductive bias. Furthermore, the method was not designed for general purpose graph learning as it relies on helper mechanisms such as neighbour sampling and a heterogeneous information network.
Recent trends in learning on graphs are dominated by architectures based on standard self-attention layers as the only learning mechanism, with a significant amount of effort put into representing the graph structure exclusively through positional and structural encodings. Graphormer [ying2021do] is one of the most prominent approaches from this class. It comes with an involved and computation-heavy suite of pre-processing steps, involving a centrality, spatial, and edge encoding. For instance, spatial encodings rely on the Floyd–Warshall algorithm that has cubic time complexity in the number of nodes and quadratic memory complexity. Also, the model employs a virtual mean-readout node that is connected to all other nodes in the graph. While Graphormer has originally been evaluated only on four datasets, the results were promising relative to GNNs and SAN. Another related approach is the Tokenized Graph Transformer (TokenGT) [kim2022pure], which treats all the nodes and edges as independent tokens. To adapt sequence learning to the graph domain, TokenGT encodes the graph information using node identifiers derived from orthogonal random features or Laplacian eigenvectors, and learnable type identifiers for nodes and edges. TokenGT is provably more expressive than standard GNNs, and can approximate $k$-WL tests with the appropriate architecture (number of layers, adequate pooling, etc.). However, this theoretical guarantee holds only with the use of positional encodings that typically break the permutation invariance over nodes that is required for consistent predictions over the same graph presented with a different node order. The main strength of TokenGT is its theoretical expressiveness, as it has only been evaluated on a single dataset where it did not outperform Graphormer.
A route involving a hybrid between classical transformers and message passing has been pursued in the GraphGPS framework [rampasek2022GPS], which combines message passing and transformer layers. As in previous works, GraphGPS puts a large emphasis on different types of encodings, proposing and analysing positional and structural encodings, further divided into local, global, and relative encodings. Exphormer [shirzad2023exphormer] is an evolution of GraphGPS that adds virtual global nodes and sparse attention based on expander graphs. While effective, such frameworks do still rely on message passing and are thus not purely attention-based solutions. Limitations include dependence on approximations (Performer [choromanski2021rethinking] for both, expander graphs for Exphormer), and decreased performance when encodings (GraphGPS) or special nodes (Exphormer) are removed. Notably, Exphormer is the first approach from this class to consider custom attention patterns given by node neighbourhoods.
@@ 2 Related Work
Figure 1: A high level overview of the GAT message passing algorithm [velickovic2018graph]. Panel A: A GAT layer receives node representations and the adjacency matrix as inputs. First a projection matrix is applied to all the nodes, which are then concatenated pairwise and passed through another linear layer, followed by a non-linearity. The final attention score is computed by the softmax function. Panel B: A GAT model stacks multiple GAT layers and uses a readout function over nodes to generate a graph-level representation. Residual connections are also illustrated as a modern enhancement of GNNs (not included in the original approach). GAT vs ESA: Our masked self-attention module relies on masking within classical scaled dot product attention which comes with different projection matrices for keys, queries, and values. Additionally, we wrap the masked scaled dot product attention with layer normalization and skip connections, the output of which is passed through an MLP. The latter is not done classically in GAT as part of its attention modules nor in the overall architecture. In contrast to GAT that operates over nodes, our masking operator is defined over sets of edges that are connected if they share a node in common. When it comes to the readout, GAT uses sum, mean or max and aggregates over nodes whereas our architecture relies on pooling by multi-head attention and aggregation over learned representations of seed vectors inherent to that module. In terms of the encoder that is responsible for learning representations of set items, ours interleaves masked and self-attention layers vertically whereas in GATs the stacked layers are entirely based on neighbourhood attention.
Figure 2: A high level overview of the SAN architecture [kreuzer2021rethinking]. Panel A: Two independent attention modules with a tied/shared value projection matrix are used for the input graph and its complement. The outputs of these two modules are convexly combined using a hyperparameter $\gamma$ before being residually added to the input. Panel B: The overall SAN architecture that stacks multiple SAN layers without residual connections, and uses a standard readout function over nodes. SAN vs ESA: What is truly different in relation to this architecture is horizontal versus vertical combination of masked and self-attention. While it is possible to set $\gamma=0$ for some layers and $\gamma=1$ for others, this avenue has never been explored in SAN and it would still result in a slightly different encoder as the query and key projections are different for the input graph and its complement. The empirical analysis in [kreuzer2021rethinking] also provides no insights relative to vertical combination of masked and self-attention, nor the impact of attention pooling which is a differentiating factor and an important part of our architecture.
@@
Figure 3: A high-level overview of Edge Set Attention and its main building blocks. Panel A: An illustration of masked scaled dot product attention with edge inputs. Edge features derived from the edge set are processed by query, key, and value projections as in standard attention. An additive edge mask, derived as in Section 2, is applied prior to the softmax operator. A mask value of 0 leaves the attention score unchanged, while a large negative value completely discards that attention score. Panel B: The Pooling by Multi-Head Attention Block (PMA). A learnable set of seed tensors $\mathbf{S}_{k}$ is randomly initialised and used as the query for a cross attention operation with the inputs. The result is further processed by Self Attention Blocks. Panel C: The Self Attention Block (SAB), illustrated here with a pre-layer-normalization architecture. The input is normalised, processed by self attention, then further normalised and processed by an MLP. The block uses residual connections. Panel D: Illustration of the Masked (Self) Attention Block (MAB). The only difference compared to SAB is the edge mask, which follows the computation outlined in Panel A. Panel E: An instantiation of the ESA model. Edge inputs are processed by various different attention blocks, here in the order MSMSP, corresponding to masked, self, masked, self, and pooling attention layers.
@@
[t]    Edge masking in PyTorch Geometric (T is the transpose; helper functions are explained in SI 6).

### v3

# An end-to-end attention-based approach for learning on graphs

@@ Abstract
There has been a recent surge in transformer-based architectures for learning on graphs, mainly motivated by attention as an effective learning mechanism and the desire to supersede the hand-crafted operators characteristic of message passing schemes. However, concerns over their empirical effectiveness, scalability, and complexity of the pre-processing steps have been raised, especially in relation to much simpler graph neural networks that typically perform on par with them across a wide range of benchmarks. To address these shortcomings, we consider graphs as sets of edges and propose a purely attention-based approach consisting of an encoder and an attention pooling mechanism. The encoder vertically interleaves masked and vanilla self-attention modules to learn an effective representation of edges while allowing for tackling possible misspecifications in input graphs. Despite its simplicity, the approach outperforms fine-tuned message passing baselines and recently proposed transformer-based methods on more than 70 node and graph-level tasks, including challenging long-range benchmarks. Moreover, we demonstrate state-of-the-art performance across different tasks, ranging from molecular to vision graphs, and heterophilous node classification. The approach also outperforms graph neural networks and transformers in transfer learning settings and scales much better than alternatives with a similar performance level or expressive power.
Introduction
@@
We empirically investigate the potential of a purely attention-based approach to learn effective representations of graph-structured data. Typically, learning on graphs is modelled as message passing, an iterative process that relies on a message function to aggregate information from a given node’s neighbourhood and an update function to incorporate the encoded message into the output representation of the node. The resulting graph neural networks (GNNs) typically stack multiple such layers to learn node representations based on vertex-rooted subtrees, essentially mimicking the one-dimensional Weisfeiler–Lehman ($1$-WL) graph isomorphism test [wl79, wlneural24]. Variations of message passing have been applied effectively in different fields such as life sciences [STOKES2020688, Wong2023, gasteiger_dimenet_2020, Merchant2023, doi:10.1126/science.abj8754, 10.1093/bioinformatics/btab804, doi:10.1126/science.add2187], electrical engineering [Chien2024], and weather prediction [doi:10.1126/science.adi2336].
@@
Despite the overall success and wide adoption of graph neural networks, several practical challenges have been identified over time. Although the message passing framework is highly flexible, the design of new layers is a challenging research problem where improvements take years to achieve and often rely on hand-crafted operators. This is particularly the case for general-purpose graph neural networks that do not exploit additional input modalities, such as atomic coordinates. For example, principal neighbourhood aggregation (PNA) is regarded as one of the most powerful message passing layers [10.5555/3495724.3496836], but it is built using a collection of manually selected neighbourhood aggregation functions, requires a degree histogram of the dataset which must be precomputed prior to learning, and further uses manually selected degree scaling. The nature of message passing also imposes certain limitations that have shaped the majority of the literature. One of the most prominent examples is the readout function used to combine node-level features into a single graph-level representation, and which is required to be permutation invariant with respect to the node order. Thus, the default choice for graph neural networks and even graph transformers remains a simple, non-learnable function such as sum, mean, or max [buterez2022graph, Buterez2023, D3SC00841J]. The limitations of this approach have been identified by Wagstaff et al. [pmlr-v97-wagstaff19a], who have shown that simple readout functions might require complex item embedding functions that are difficult to learn using standard neural networks. Additionally, graph neural networks have shown limitations in terms of over-smoothing [gcnsmooth, digiovanni2023understanding, sheaf22], linked to node representations becoming similar with increased depth, and over-squashing [alon2021on, topping2022understanding] due to information compression through bottleneck edges. The former has been associated with poor performance on node classification tasks with heterophilic graphs and it is hypothesized that this is due to GNNs acting as low-pass filters. Recently, Di Giovanni et al. [digiovanni2023understanding] have studied over-smoothing using gradient flows on graphs and have demonstrated that some time-continuous GNNs are indeed dominated by low frequencies. Moreover, behaviour opposite to over-smoothing known as over-sharpening has been identified in a setting with linear graph convolutions and symmetric weights, achieved via repulsion induced by negative eigenvalues of the weight matrix. Over-squashing, on the other hand, impacts performance in cases where information from distant nodes is relevant for predictive tasks and it has been linked to graph bottlenecks, characterized by a rapid increase in the number of $k$-hop neighbourhoods as $k$ or the number of layers increases. Topping et al. [topping2022understanding] have provided a theoretical characterization of over-squashing and introduced a notion of graph curvature as means for quantifying it, along with a graph rewiring algorithm known as stochastic discrete Ricci flow that can mitigate some of the bottleneck effects. Alternative solutions for these two problems typically take the form of message regularisation schemes [godwin2022simple, Zhao2020PairNorm, cai2021graphnorm]. There is, however, no clear consensus on the right architectural choices for building effective deep message passing neural networks and effectively addressing all of those challenges. Transfer learning and strategies such as pre-training and fine-tuning are also less ubiquitous in graph neural networks because of modest or ambiguous benefits, as opposed to large language models [Hu2020Strategies].
@@
Figure 1: A – B. A high level overview of the GAT message passing algorithm [velickovic2018graph]. A. A GAT layer receives node representations and the adjacency matrix as inputs. First a projection matrix is applied to all the nodes, which are then concatenated pairwise and passed through another linear layer, followed by a non-linearity. The final attention score is computed by the softmax function. B. A GAT model stacks multiple GAT layers and uses a readout function over nodes to generate a graph-level representation. Residual connections are also illustrated as a modern enhancement of GNNs (not included in the original approach). C – D. A high level overview of the SAN architecture [kreuzer2021rethinking]. C. Two independent attention modules with a shared value projection matrix are used for the input graph and its complement. The outputs of these two modules are convexly combined using a hyperparameter $\gamma$ before being residually added to the input. D. The overall SAN architecture that stacks multiple SAN layers without residual connections, and uses a standard readout function over nodes. The differences between ESA, GAT, and SAN are detailed in SI 1.
The attention mechanism [NIPS2017_3f5ee243] is one of the main sources of innovation within graph learning, either by directly incorporating attention within message passing [velickovic2018graph, brody2022how], by formulating graph learning as a language processing task [ying2021do, kim2022pure], or by combining vanilla GNN layers with attention layers [rampasek2022recipe, shirzad2023exphormer, buterez2022graph, Buterez2023, deng2024polynormer]. However, several concerns have been raised regarding the performance, scalability, and complexity of such methods. Performance-wise, recent reports indicate that sophisticated graph transformers underperform compared to simple but tuned GNNs [tonshoff2023where, luo2024classic]. This line of work highlights the importance of empirically evaluating new methods relative to strong baselines. Separately, recent graph transformers have focused on increasingly more complex helper mechanisms, such as computationally expensive preprocessing and learning steps [ying2021do], various different encodings (e.g. positional, structural, and relational) [ying2021do, kim2022pure, rampasek2022recipe, 10.5555/3618408.3619379], inclusion of virtual nodes and edges [ying2021do, shirzad2023exphormer], conversion of the problem to natural language processing [ying2021do, kim2022pure], and other non-trivial graph transformations [shirzad2023exphormer, 10.5555/3618408.3619379]. These complications can significantly increase computational requirements, reducing the chance of being widely adopted and replacing GNNs.
Motivated by the effectiveness of attention as a learning mechanism and recent advances in efficient and exact attention, we introduce an end-to-end attention-based architecture for learning on graphs that is simple to implement, scalable, and achieves state-of-the-art results. The proposed architecture considers graphs as sets of edges, leveraging an encoder that interleaves masked and self-attention mechanisms to learn effective representations. The attention-based pooling component mimics the functionality of a readout function and is responsible for aggregating the edge-level features into a permutation-invariant graph-level representation. The masked attention mechanism allows for learning effective edge representations originating from the graph connectivity, and the combination with self-attention layers vertically allows for expanding on this information while having a strong prior. Masking can thus be seen as leveraging specified relational information and its vertical combination with self-attention as a means to overcome possible misspecification of the input graph. The masking operator is injected into the pairwise attention weight matrix and allows only for attention between linked primitives. For a pair of edges, connectivity translates to having a shared node between them. We focus primarily on learning through edge sets due to empirically high performance and refer to our architecture as edge-set attention (ESA). We also show that the overall architecture is effective in propagating information across nodes through dedicated node-level benchmarks. The ESA architecture is general purpose, in the sense that it relies only on the graph structure and possibly node and edge features, and it is not restricted to any particular domain. Furthermore, ESA does not rely on positional, structural, relative, or similar encodings, it does not encode graph structures as tokens or other language (sequence) specific concepts, and it does not require any pre-computations.
@@
Figure 2: A high-level overview of Edge Set Attention and its main building blocks. A. An illustration of masked scaled dot product attention with edge inputs. Edge features derived from the edge set are processed by query, key, and value projections as in standard attention. An additive edge mask, derived as in Introduction, is applied prior to the softmax operator. A mask value of 0 leaves the attention score unchanged, while a large negative value completely discards that attention score. B. The Pooling by Multi-Head Attention Block (PMA). A learnable set of seed tensors $\mathbf{S}_{k}$ is randomly initialised and used as the query for a cross attention operation with the inputs. The result is further processed by Self Attention Blocks. C. The Self Attention Block (SAB), illustrated here with a pre-layer-normalization architecture. The input is normalised, processed by self attention, then further normalised and processed by an MLP. The block uses residual connections. D. Illustration of the Masked (Self) Attention Block (MAB). The only difference compared to SAB is the edge mask, which follows the computation outlined in panel A. E. An instantiation of the ESA model. Edge inputs are processed by various different attention blocks, here in the order MSMSP, corresponding to masked, self, masked, self, and pooling attention layers.
Despite its apparent simplicity, ESA-based learning overwhelmingly outperforms strong and tuned GNN baselines and much more involved transformer-based models. Our evaluation is extensive, totalling $70$ datasets and benchmarks from different domains such as quantum mechanics, molecular docking, physical chemistry, biophysics, bioinformatics, computer vision, social networks, functional call graphs, and synthetic graphs. At the node level, we include both homophilous and heterophilous graph tasks, as well as shortest path problems, as these require modelling long-range interactions. Beyond supervised learning tasks, we explore the potential for transfer learning in the context of drug discovery and quantum mechanics [Buterez_Janet_Kiddle_Oglic_Liò_2024, doi:10.1021/acs.jcim.2c01569] and show that ESA is a viable transfer learning strategy compared to vanilla GNNs and graph transformers.
Related work. An attention mechanism that mimics message passing and that limits the attention computations to neighbouring nodes has been first proposed in GAT [velickovic2018graph]. We consider masking as an abstraction of the GAT attention operator that allows for building general-purpose relational structures between items in a set (e.g. $k$-hop neighbourhoods or conformational masks extracted from $3$D molecular structures). The attention mechanism in GAT is implemented as a single linear projection matrix that does not explicitly distinguish between keys, queries, and values as in standard dot product attention and an additional linear layer after concatenating the representations of connected nodes, along with a non-linearity. This type of simplified attention has been labelled in subsequent work as static and was shown to have limited expressive power [brody2022how]. Brody et al. instead proposed dynamic attention, a simple reordering of operations in GAT, resulting in a more expressive GATv2 model [brody2022how] – however, at the price of doubling the parameter count and the corresponding memory consumption. A high-level overview of GAT in the context of masked attention is provided in Figure 1A and B, along with the main differences to the proposed architecture. Similarly to GAT, several adaptations of the original scaled dot product attention have been proposed for graphs [ijcai2021p214, dwivedi2021generalization], where the focus was on defining an attention mechanism constrained by node connectivity and replacing the positional encodings of the original transformer model with more appropriate graph alternatives, such as Laplacian eigenvectors. These approaches, while interesting and forward-looking, did not convincingly outperform simple GNNs [ijcai2021p214, dwivedi2021generalization]. Building on this line of work, an architecture that can be seen as an instance of masked transformers has been proposed in SAN [kreuzer2021rethinking], illustrated in Figure 1C and D. The attention coefficients are defined as a convex combination of the scores (controlled by hyperparameter $\gamma$) associated with the original graph and its complement. Min et al. [10.1145/3477495.3532031] have also considered a masking mechanism for standard transformers. However, the graph structure itself is not used directly in the masking process as the devised masks correspond to four types of prior interaction graphs (induced, similarity, cross neighbourhood, and complete subgraphs), acting as an inductive bias. Furthermore, the method was not designed for general purpose graph learning as it relies on helper mechanisms such as neighbour sampling and a heterogeneous information network.
Recent trends in learning on graphs are dominated by architectures based on standard self-attention layers as the only learning mechanism, with a significant amount of effort put into representing the graph structure exclusively through positional and structural encodings. Graphormer [ying2021do] is one of the most prominent approaches of this class. It comes with an involved and computation-heavy suite of pre-processing steps, involving a centrality, spatial, and edge encoding. For instance, spatial encodings rely on the Floyd–Warshall algorithm, which has cubic time complexity in the number of nodes and quadratic memory complexity. In addition, the model employs a virtual mean readout node that is connected to all other nodes in the graph. Although Graphormer was originally evaluated on only four datasets, the results were promising relative to GNNs and SAN. Another related approach is the Tokenized Graph Transformer (TokenGT) [kim2022pure], which treats all nodes and edges as independent tokens. To adapt sequence learning to the graph domain, TokenGT encodes the graph information using node identifiers derived from orthogonal random features or Laplacian eigenvectors, and learnable type identifiers for nodes and edges. TokenGT is provably more expressive than standard GNNs and can approximate $k$-WL tests with the appropriate architecture (number of layers, adequate pooling, etc.). However, this theoretical guarantee holds only with the use of positional encodings that typically break the permutation invariance over nodes that is required for consistent predictions over the same graph presented with a different node order. The main strength of TokenGT is its theoretical expressiveness, as it has only been evaluated on a single dataset where it did not outperform Graphormer.
A route involving a hybrid between classical transformers and message passing has been pursued in the GraphGPS framework [rampasek2022GPS], which combines message passing and transformer layers. As in previous works, GraphGPS places a large emphasis on different types of encodings, proposing and analysing positional and structural encodings, further divided into local, global, and relative encodings. Exphormer [shirzad2023exphormer] is an evolution of GraphGPS that adds virtual global nodes and sparse attention based on expander graphs. While effective, such frameworks still rely on message passing and are thus not purely attention-based solutions. Limitations include dependence on approximations (Performer [choromanski2021rethinking] for both approaches, expander graphs for Exphormer), and decreased performance when encodings (GraphGPS) or special nodes (Exphormer) are removed. Notably, Exphormer is the first approach from this class to consider custom attention patterns given by node neighbourhoods. Polynormer [deng2024polynormer] is another approach from this class that combines local (mimicking graph convolutions or GATs) and global attention (scaled dot product attention), and achieves polynomial expressivity by a gating mechanism which can be seen as an alternative to residual connections, with element-wise multiplication between the projected inputs and outputs of the attention blocks.
@@
[t]    Edge masking in PyTorch Geometric (T is the transpose; helper functions are explained in SI 8).